A wire of length 0.75 m carries a current of 20.0 A. The wire experiences a force of 5.8 N in the presence of a magnetic field of 0.50 T. What is the angle between the magnetic field and the wire?

1
Expert's answer
2021-10-31T18:15:33-0400
"F_B=BILsin\theta,""\theta=sin^{-1}(\dfrac{F_B}{BIL})=sin^{-1}(\dfrac{5.8\ N}{0.50\ T\times20\ A\times0.75\ m})=50.6^{\circ}"